Two adjacent energy levels of an electron in a harmonic potential well are known to be 1.10 eV and 1.54 eV .

What is the spring constant of the potential well?
Answer #1

En = ( n+1/2)ħω

adjacent energy levels difference = En+1 -En = ħω

ħω = 1.54-1.10 = 0.44 ev

ω = 0.44/ħ = 0.44/ 6.5821 × 10-16eV s

=6.68479665e14

ω = sqrt(k/m)

k = m*ω2 = 9.1*10^-31*(6.68479665e14 )^2 = 0.4066472 N/m  

spring constant of the potential well =  0.4066472 N/m  

****************************************************************************
Goodluck for exam Comment in case any doubt, will reply for sure..